An 90.0 kg spacewalking astronaut pushes off a 660 kg satellite, exerting a 85.0 N force for the 0.590 s it takes him to straighten his arms.
Answer #1

given
mass of astronaut, m = 90 kg
mass of satellite, M = 660 kg
force exerted by astronaut on the satellite, F = 85 N
time dueing which this force is applied, t = 0.59 s
hence
acceleration of the satellite = F/M = 0.12878 m/s/s
t = 0.59 s
hence
final speed of the satellite, v = Ft/M = 0.0759848 m/s

from ocnservation of moemntum
final speed of the astronaut = u
mu = Mv
u = (Mv/m) = 0.55722 m/s